Solve;Β Β (𝑧 2 βˆ’ 2𝑦𝑧 βˆ’ 𝑦 2 )𝑝 + (π‘₯𝑦 + 𝑧π‘₯)π‘ž = π‘₯𝑦 βˆ’ 𝑧x


1
Expert's answer
2021-06-14T11:14:47-0400

"(z^2-2yz-y^2)z_x+(xy+zx)z_y= xy-zx\\
\text{By Charpit-Lagrange method :}\\
\frac{dx}{z2−2yz−y2}=\frac{dy}{xy+zx}=\frac{dz}{xy−zx}\\
\text{First characteristic equation,}\\ \frac{
dy}{xy+zx}=\frac{
dz}{xy−zx}\\\frac{
dz}{dy}=\frac{y−z}{y+z} \\
z^2+2yz−y^2=c1\\
\text{Second characteristic equation,}\\
\frac{xdx+ydy+zdz}{x(z^2-2yz-y^2)+y(xy+zx)+z(xy-zx)}=0\\\frac{xdx+ydy+zdz}{0}=0\\\implies xdx+ydy+zdz=0\\
x^2+y^2+z^2=c2
\\
\text{Therefore, solution is given by,}\\
x^2+y^2+z^2=f(z^2+2yz-y^2)"